OpenStudy (anonymous):

How do I simplify [(-x^5)]6? The brackets are confusing me.

OpenStudy (anonymous):

\([(-x^5)]6 = [(-x^5)]\times 6\)

Parth (parthkohli):

Maybe they meant \((-x^5)^6\).

OpenStudy (anonymous):

Did you get that @JamesR4494 ?

OpenStudy (anonymous):

If so , then @JamesR4494 please tell.

Parth (parthkohli):

In that case, try using \((-x^5) \times (-x^5) \times (-x^5) \times (-x ^5 ) \times (-x^5) \times (-x^5)\)

Parth (parthkohli):

You can make that even better by doing this:\[(-1 \times x^5) \times (-1 \times x^5) \times (-1 \times x^5) \times (-1 \times x^5) \times (-1 \times x^5) \times (-1 \times x^5)\]

OpenStudy (anonymous):

or : \((-x^5)^6 = (-x^{5\times 6}) = (-x^{30})\)

Parth (parthkohli):

@SheldonEinstein This isn't so simple, man :-)

OpenStudy (anonymous):

I meant \(x^{30}\)

Parth (parthkohli):

Remember:\[(-1 \times x^5) ^6 = (-1)^6 \times (x^5)^6\]

OpenStudy (anonymous):

\((-x^{5})^6 = (x^{5\times 6} ) = x^{30}\) By using the property \((a^m)^n = a^{mn}\)

Parth (parthkohli):

Yup, but you also have to explain why the minus disappeared :-)